OpenStudy (anonymous):

A rectangle has an area of 88 square units. The width of the rectangle is 8 units. The length of the rectangle is 4x - 1. What is the measure of the rectangle's length? 2 3 7 11

OpenStudy (ajprincess):

 Area=Length*Width 88=(4x-1)*8 4x-1=88/8 4x-1=11 Add 1 both sides and then divide by 4 to find x.